# Noncollinear Spin-Flip TDDFT for Potential Energy Surface Crossings: Conical Intersections and Spin Crossings

### v2

@@ 3.2 Avoiding Spurious PES Crossings
In the case of pure exchange-correlation functionals, the excitation energies obtained from collinear spin-flip TDDFT are simply the energy differences between the occupied spin-up and virtual spin-down molecular orbitals. When the two singly-occupied orbitals (the O orbitals in Figure 1) are (near-)degenerate, either due to symmetry-imposed degeneracy or accidental degeneracy, the calculated excitation energies can become (near-)degenerate as well, resulting in physically incorrect state degeneracies.
@@
Figure 3 illustrates this problem using two examples. Figure 3a shows the potential energy curves for the ground state and three low-lying excited states of ethylene as a function of the torsion angle, calculated with collinear spin-flip TDDFT (BLYP/cc-pVDZ63, 64, 65). At the 90°90  ° twisted geometry, the $\mathrm{D}_{2d}$ point group imposes an exact degeneracy of frontier orbitals, which in turn enforces an artificial fourfold degeneracy predicted by collinear spin-flip TDDFT. A similar issue is observed for stilbene (Figure 3c), where an accidental near-degeneracy of the frontier orbitals arises at 90°90  ° torsion ($\mathrm{C}_{2}$ symmetry, calculated with BLYP/6-31G56, 57). As pointed out by Krylov and co-workers34, introducing hybrid functionals into collinear spin-flip TDDFT can alleviate the issue to some extent by lifting finite energy gaps. However, the magnitude of these gaps is highly sensitive to the fraction of HF exchange in the functional, making the remedy somewhat ad hoc.
@@
Figure 3: Potential energy curves for ethylene (a, b) and stilbene (c, d) as a function of torsion angle, computed using collinear (a, c) and noncollinear (b, d) spin-flip TDDFT with BLYP functional. SC stands for strongly spin-contaminated states with $\langle{\hat{S}^{2}}\rangle\approx 1$. Geometries of twisted stilbene are generated using Molclus66.
@@
In contrast, noncollinear spin-flip TDDFT naturally lifts these degeneracies even when pure functionals are used, as shown in Figure 3b and 3d. This demonstrates an advantage of the noncollinear approach: it inherently avoids spurious state crossings that arise from orbital degeneracies, thereby providing a more robust description of the order of electronic states. It is also worth Figure 3b and 3d closely resemble the TDDFT-1D results reported by Teh and Subotnik,29 where doubly excited configurations are explicitly included, highlighting the importance of double excitations. In the case of noncollinear spin-flip TDDFT, by contrast, these double excitations are incorporated naturally within the framework, without the need for explicit inclusion.
Finally, we point out an interesting phenomenon: at the 90°90  ° twisted geometry of ethylene and stilbene, the noncollinear spin-flip TDDFT results reveal near-degeneracies between $\mathrm{S}_{1}$ and $\mathrm{S}_{2}$, as well as between $\mathrm{S}_{0}$ and $\mathrm{T}$. This kind of near-degeneracy is a common phenomenon among diradicals67, 68.
@@ 3.3 Locating $\mathrm{S}_{0}/\mathrm{T}$ Crossings
Noncollinear spin-flip TDDFT has the important property of preserving the degeneracy between the high-spin triplet reference state $|{\mathrm{T}(S_{z}=1)}\rangle$ and the spin-flip triplet state $|{\mathrm{T}(S_{z}=0)}\rangle$50, as shown in Figure 1. This feature, which is absent in collinear spin-flip TDDFT, facilitates the use of noncollinear spin-flip TDDFT for locating singlet-triplet intersections.
@@
Figure 4: (a) MECP geometries of nitroxyl calculated by different methods. Structures are visualized using VMD69, 70. (b) PESs of $\mathrm{S}_{0}$ and $\mathrm{T}$ states near the crossing point calculated by noncollinear spin-flip TDDFT. BLA stands for bond length alternation.
@@
[…]
@@
To verify that noncollinear TDDFT can provide the correct topography near the $\mathrm{S}_{0}/\mathrm{T}$ crossing point, we present the PESs near the crossing point in Figure 4b. It can be seen that the degeneracy can only be lifted along a single direction, resulting in a line-shaped crossing seam. This stands in sharp contrast to the $\mathrm{S}_{0}/\mathrm{S}_{1}$ conical intersection, where two effective directions lift the degeneracy.
@@ 3.4 Locating $\mathrm{S}_{1}/\mathrm{T}$ Crossings
Aside from the $\mathrm{S}_{0}/\mathrm{T}$ crossings, noncollinear spin-flip TDDFT can also be used to locate the $\mathrm{S}_{1}/\mathrm{T}$ crossings. As an example, $\mathrm{S}_{1}/\mathrm{T}$ crossing of $o$-nitrophenol is located at B3LYP/6-31G** level, with the structure shown in Figure 5. It shows that the geometry is quite similar to the reference results calculated by 6SA-CASSCF(10,10)/6-31G** in Ref. 71.
@@
It should be noted that locating the crossing point between $\mathrm{S}_{1}$ and $\mathrm{T}$ is not trivial. In particular, it requires the use of electronic structure methods that allow the energy of $\mathrm{S}_{1}$ to be lower than that of $\mathrm{T}$, which represents a violation of Hund’s rule. Conventional spin-conserving TDDFT typically fails to achieve this72, whereas spin-flip TDDFT can73, 74.
@@ 4 Assessment of the Potential of Noncollinear Spin-Flip TDDFT for Nonadiabatic Dynamics Simulations
    Description of multiple PESs with correct spin multiplicities. Nonadiabatic dynamics relies on multiple potential energy surfaces to describe the evolution of a system, making this a prerequisite condition. Noncollinear spin-flip TDDFT naturally describes several low-lying singlet and triplet states, making it well suited for simulating nonadiabatic couplings among these states. Meanwhile, compared to collinear spin-flip TDDFT, spin contamination is significantly reduced. However, to fully eliminate spin contamination, especially for higher excited states, spin-adaptation techniques are still required75, 76, 77, 78, 79, 80.
@@
    Analytic gradients of PESs. Efficient nonadiabatic dynamics requires analytic gradients of PESs to propagate nuclear wavefunctions. With the numerically stable spin-flip kernel (Eq. 7) and its derivatives, noncollinear spin-flip TDDFT can provide analytic gradients49.
@@
    Correct ordering of states. An roughly correct ordering of electronic states is important because it determines whether potential energy surface crossings exist within a given range of molecular geometries. We have shown an example in Section 3.4 where, unlike spin-conserving TDDFT that always predicts the energy of $\mathrm{S}_{1}$ to be higher than that of $\mathrm{T}$, spin-flip TDDFT can meaningfully predict whether $\mathrm{S}_{1}$ lies below $\mathrm{T}$. Molecules with small or even negative $\mathrm{S}_{1}/\mathrm{T}$ gaps at their ground-state geometries, such as thermally activated delayed fluorescence (TADF) molecules, are of great significance in organic light-emitting diodes (OLEDs)72, 73, 81, 82. The two molecules shown in Figure 6, cyclazine and heptazine, are typical examples believed to have $\mathrm{S}_{1}$ energies lower than $\mathrm{T}$. To investigate their energetic behavior, we optimize the geometries at restricted Kohn-Sham (RKS) B3LYP/6-31G* level using Gaussian 1683. Vertical excitation energies are then computed using both spin-conserving and spin-flip TDDFT. The results in Table 3 clearly show that noncollinear spin-flip TDDFT (except for the SVWN functional) predicts the correct ordering of $\mathrm{S}_{1}$ and $\mathrm{T}$ states, while spin-conserving TDDFT always predicts $\mathrm{S}_{1}$ to lie above $\mathrm{T}$. Finally, it is worth mentioning that degeneracy of potential energy surfaces can also be regarded as a special case of relative ordering. The degeneracy between the $|{\mathrm{T}(S_{z}=1)}\rangle$ and $|{\mathrm{T}(S_{z}=0)}\rangle$ is strictly preserved in noncollinear spin-flip TDDFT50.
@@
    Table 3: Singlet-triplet gap $E(\mathrm{S}_{1})-E(\mathrm{T})$ (eV) computed with different functionals.
@@
    Cyclazine SVWN 0.149 0.039 0.131-0.131- 0.131
    PBE 0.176 0.034-0.034- 0.034
    TPSS 0.189 0.100-0.100- 0.100
    B3LYP 0.202 0.281-0.281- 0.281
    Heptazine SVWN 0.155 0.023 0.219-0.219- 0.219
    PBE 0.178 0.054-0.054- 0.054
    TPSS 0.189 0.120-0.120- 0.120
    B3LYP 0.210 0.353-0.353- 0.353
    aTheoretically best estimate values from Ref. 84.
@@
    Correct topology at potential energy surface crossings. Regions where two potential energy surfaces intersect are where nonadiabatic processes efficiently take place. Therefore, providing a topologically correct description of the potential energy surfaces near such crossing points is a critical factor for achieving reliable nonadiabatic dynamics. Noncollinear spin-flip TDDFT meets this requirement. As demonstrated in Sections 3.1 and 3.3, it can accurately describe the conical intersections between singlet states as well as the crossing seams between singlet and triplet states.
@@
    Nonadiabatic derivative coupling and spin-orbit coupling. They are the driving forces behind nonadiabatic transitions between electronic states of the same and different spin multiplicities, respectively. Therefore, they are essential ingredients in nonadiabatic dynamics methods such as Tully’s fewest switches surface hopping85, 86, 87. Similar to the treatment of analytic gradients32, noncollinear spin-flip TDDFT is capable of providing analytic nonadiabatic derivative couplings, a feature that is currently under development. However, due to slight spin contamination, nonadiabatic couplings between states of different multiplicities may be nonzero, and fully avoiding such unphysical couplings relies on appropriate spin-adaptation within the noncollinear spin-flip TDDFT framework. Moreover, one notable advantage of the noncollinear spin-flip TDDFT approach is that it offers a consistent framework for treating both derivative couplings and spin-orbit coupling effects simultaneously, which is particularly beneficial for systems where intersystem crossing and internal conversion are energetically competitive88.
@@
    Capability to handle multiconfigurational character. Nonadiabatic dynamics simulations may involve significant changes in bond lengths and angles, and even chemical reactions. During such processes, the relative weights of electronic configurations can vary considerably. Therefore, electronic structure methods capable of describing bond breaking and formation are highly desirable. Our previous work has examined how noncollinear spin-flip TDDFT handles single bond dissociation49, 50, but the dissociation of double bonds remains a challenge.

### v3

@@ 3.2 Avoiding Spurious PES Crossings
For pure exchange-correlation functionals, the excitation energies obtained from collinear spin-flip TDDFT are simply the energy differences between the occupied spin-up and virtual spin-down molecular orbitals. When the two singly-occupied orbitals (the O orbitals in Figure 1) are (near-)degenerate, either due to symmetry-imposed degeneracy or accidental degeneracy, the resulting excitation energies can become (near-)degenerate as well, resulting in physically incorrect state degeneracies.
@@
Figure 3 illustrates this problem using two examples. Figure 3a shows the potential energy curves for the ground state and three low-lying excited states of ethylene as a function of the torsion angle, calculated with collinear spin-flip TDDFT (BLYP/cc-pVDZbecke1988densityfunctional, lee1988development, dunning1989gaussian). At the 90° twisted geometry, the $\mathrm{D}_{2d}$ point group imposes an exact degeneracy of frontier orbitals, which in turn enforces an artificial fourfold degeneracy predicted by collinear spin-flip TDDFT. A similar issue is observed for stilbene (Figure 3c), where an accidental near-degeneracy of the frontier orbitals arises at 90° torsion ($\mathrm{C}_{2}$ symmetry, calculated with BLYP/6-31Gditchfield1971selfconsistent, hehre1972self). Such spurious degeneracies originate from the inherent limitation of collinear spin-flip TDDFT when using pure functionals, and are independent of the choice of basis set or functional. As pointed out by Krylov and co-workersshao2003spin, introducing hybrid functionals into collinear spin-flip TDDFT can partially alleviate the issue to some extent by lifting finite energy gaps. However, the magnitude of these gaps can be arbitrarily tuned by adjusting the fraction of HF exchange in the functional, rendering the remedy somewhat ad hoc.
@@
Figure 3: Potential energy curves for ethylene (a, b) and stilbene (c, d) as a function of torsion angle, computed using collinear (a, c) and noncollinear (b, d) spin-flip TDDFT with BLYP functional. SC stands for strongly spin-contaminated states with $\langle{\hat{S}^{2}}\rangle\approx 1$. Geometries of twisted stilbene are generated using Molclusmolclus.
@@
In contrast, noncollinear spin-flip TDDFT naturally lifts these degeneracies even when pure functionals are used, as shown in Figure 3b and 3d. This demonstrates an advantage of the noncollinear approach: it inherently avoids spurious state crossings that arise from orbital degeneracies, thereby providing a more robust description of the order of electronic states. It is also worth noting that Figure 3b and 3d closely resemble the TDDFT-1D results reported by Teh and Subotnik,teh2019simplest where doubly excited configurations are explicitly included, highlighting the importance of double excitations. Such doubly-excited configuration $|{\mathrm{DES}}\rangle$ is incorporated naturally within noncollinear spin-flip TDDFT. As a direct demonstration of this, Appendix B presents several examples of excited states dominated by $|{\mathrm{DES}}\rangle$.
It is worth noting that, at the 90° twisted geometry of ethylene and stilbene, noncollinear spin-flip TDDFT predicts near-degeneracies between $\mathrm{S}_{1}$ and $\mathrm{S}_{2}$, as well as between $\mathrm{S}_{0}$ and $\mathrm{T}$. Such near-degeneracies are well-acknowledged features of diradical systemspittner1999assessment, stuyver2019diradicals.
@@ 3.3 Locating $\mathrm{S}_{0}/\mathrm{T}$ Crossings
A key requirement for locating intersections between singlet and triplet states is that the spin-flip TDDFT formalism should yield a unique and consistent energy for all components of a given triplet state. In practice, however, collinear spin-flip TDDFT breaks this degeneracy, producing an artificial energy splitting of roughly 1 eV between the reference state $|{\mathrm{T}(S_{z}=1)}\rangle$ and the spin-flip state $|{\mathrm{T}(S_{z}=0)}\rangle$. This self-splitting problem introduces ambiguity in defining the triplet energy and consequently affects the identification of singlet–triplet crossings.
In contrast, noncollinear spin-flip TDDFT rigorously preserves the degeneracy between different $S_{z}$ components of the same spin multiplet, as formally proven in Ref. wang2025zero and illustrated in Figure 1. It is worth noting that this theorem further holds for any high-spin reference with nonzero $S_{z}$ (e.g., $S_{z}=\tfrac{1}{2},1,\tfrac{3}{2},\dots$).
@@
Figure 4: (a) MECP geometries of nitroxyl calculated by different methods. (b) PESs of $\mathrm{S}_{0}$ and $\mathrm{T}$ states near the crossing point calculated by noncollinear spin-flip TDDFT. BLA stands for bond length alternation.
@@
Methods RR(NH) ($\mathrm{\SIUnitSymbolAngstrom}$) RR(NO) ($\mathrm{\SIUnitSymbolAngstrom}$) $\angle$(HNO) ($\mathrm{\SIUnitSymbolDegree}$) $\Delta E$ (eV)
@@
To verify that noncollinear TDDFT can provide the correct topography near the $\mathrm{S}_{0}/\mathrm{T}$ crossing point, we present the PESs near the crossing point in Figure 4b. It can be seen that the degeneracy is only lifted along a single direction, resulting in a line-shaped crossing seam. This stands in sharp contrast to the $\mathrm{S}_{0}/\mathrm{S}_{1}$ conical intersection, where two effective directions lift the degeneracy.
@@ 3.4 Locating $\mathrm{S}_{1}/\mathrm{T}$ Crossings
Aside from the $\mathrm{S}_{0}/\mathrm{T}$ crossings, noncollinear spin-flip TDDFT can also be used to locate the $\mathrm{S}_{1}/\mathrm{T}$ crossings. As an example, $\mathrm{S}_{1}/\mathrm{T}$ crossing of $o$-nitrophenol is located at B3LYP/6-31G** level, with the structure shown in Figure 5. It shows that the geometry is quite similar to the reference results calculated by 6SA-CASSCF(10,10)/6-31G** in Ref. xu2015photoisomerization.
Figure 5: (a) Chemical structure of $o$-nitrophenol. (b) Geometry of the $\mathrm{S}_{1}/\mathrm{T}$ crossing point calculated by noncollinear spin-flip TDDFT. The numbers indicate the bond lengths (in $\mathrm{\SIUnitSymbolAngstrom}$) and dihedral angles (in $\mathrm{\SIUnitSymbolDegree}$), with the values in parentheses corresponding to the reference results.
It should be noted that locating the crossing point between $\mathrm{S}_{1}$ and $\mathrm{T}$ is not trivial. In particular, it requires the use of electronic structure methods that allow the energy of $\mathrm{S}_{1}$ to be lower than that of $\mathrm{T}$, which represents a violation of Hund’s rule. Conventional spin-conserving TDDFT typically fails to achieve thisdesilva2019inverted, whereas spin-flip TDDFT canpollice2021organic, lashkaripour2025addressing.
3.5 $\mathrm{D}_{0}/\mathrm{D}_{1}$ Conical Intersections for Odd-Electron Systems
@@
However, it is worth noting that for systems with an odd number of electrons, conical intersections (M2M-2 dimensional intersection space) only occur in the absence of spin-orbit coupling. When spin-orbit coupling is included, each potential energy surface corresponds to a pair of time-reversal-related states sharing the same energy (unless time-reversal symmetry is broken by an external magnetic field). In such cases, the intersection space dimensionality becomes M5M-5mead1979noncrossing.
[…]
@@ 4 Assessment of the Potential of Noncollinear Spin-Flip TDDFT for Nonadiabatic Dynamics Simulations
    Description of multiple PESs with correct spin multiplicities. Nonadiabatic dynamics relies on multiple potential energy surfaces to describe the evolution of a system, making this a prerequisite condition. Noncollinear spin-flip TDDFT naturally describes several low-lying singlet and triplet states, making it well suited for simulating nonadiabatic couplings among these states. Meanwhile, compared to collinear spin-flip TDDFT, spin contamination is significantly reduced. However, to fully eliminate spin contamination, especially for higher excited states, spin-adaptation techniques are still requiredli2010spinadapted, li2011spinadapted, li2011spinadapteda, zhang2015spinflip, lee2018eliminating, park2023mixedreference.
@@
    Analytic gradients of PESs. Efficient nonadiabatic dynamics requires analytic gradients of PESs to propagate nuclear wavefunctions. With the numerically stable spin-flip kernel (Eq. 7) and its derivatives, noncollinear spin-flip TDDFT can provide analytic gradientsli2025analytic.
@@
    Correct ordering of states. An roughly correct ordering of electronic states is important because it determines whether potential energy surface crossings exist within a given range of molecular geometries. We have shown an example in Section 3.4 where, unlike spin-conserving TDDFT that always predicts the energy of $\mathrm{S}_{1}$ to be higher than that of $\mathrm{T}$, spin-flip TDDFT can meaningfully predict whether $\mathrm{S}_{1}$ lies below $\mathrm{T}$. Molecules with small or even negative $\mathrm{S}_{1}/\mathrm{T}$ gaps at their ground-state geometries, such as thermally activated delayed fluorescence (TADF) molecules, are of great significance in organic light-emitting diodes (OLEDs)desilva2019inverted, pollice2021organic, won2023inverted, perez-jimenez2025role. The two molecules shown in Figure 7, cyclazine and heptazine, are typical examples believed to have $\mathrm{S}_{1}$ energies lower than $\mathrm{T}$. Vertical excitation energies of two molecules are computed using both spin-conserving and spin-flip TDDFT at the geometries reported in Ref. loos2023heptazine, where they were optimized at CCSD(T)/cc-pVTZ level. The results in Table 3 clearly show that noncollinear spin-flip TDDFT (except for the SVWN functional) predicts the correct ordering of $\mathrm{S}_{1}$ and $\mathrm{T}$ states, while spin-conserving TDDFT always predicts $\mathrm{S}_{1}$ to lie above $\mathrm{T}$. Finally, it is worth mentioning that degeneracy of potential energy surfaces can also be regarded as a special case of relative ordering. The degeneracy between the $|{\mathrm{T}(S_{z}=1)}\rangle$ and $|{\mathrm{T}(S_{z}=0)}\rangle$ is strictly preserved in noncollinear spin-flip TDDFTwang2025zero.
@@
    Table 3: Singlet-triplet gap $E(\mathrm{S}_{1})-E(\mathrm{T})$ (eV) computed with different functionals using cc-pVTZ basis set.
@@
    Cyclazine SVWN 0.147 $0.049$ 0.131-0.131
    BLYP 0.166 0.024-0.024
    B3LYP 0.205 0.240-0.240
    PBE 0.179 0.019-0.019
    PBE0 0.241 0.365-0.365
    Heptazine SVWN 0.159 $0.034$ 0.219-0.219
    BLYP 0.176 0.045-0.045
    B3LYP 0.220 0.322-0.322
    PBE 0.187 0.043-0.043
    PBE0 0.254 0.481-0.481
    aTheoretically best estimate values from Ref. loos2023heptazine.
@@
    Correct topology at potential energy surface crossings. Regions where two potential energy surfaces intersect are where nonadiabatic processes efficiently take place. Therefore, providing a topologically correct description of the potential energy surfaces near such crossing points is a critical factor for achieving reliable nonadiabatic dynamics. Noncollinear spin-flip TDDFT meets this requirement. As demonstrated in Sections 3, it correctly describes the conical intersections between singlet states as well as the crossing seams between singlet and triplet states.
@@
    Nonadiabatic derivative coupling and spin-orbit coupling. They are the driving forces behind nonadiabatic transitions between electronic states of the same and different spin multiplicities, respectively. Therefore, they are essential ingredients in nonadiabatic dynamics methods such as Tully’s fewest switches surface hoppingtully1990molecular, cui2014generalized, crespo-otero2018recent. Similar to the treatment of analytic gradientszhang2014analytic, noncollinear spin-flip TDDFT is capable of providing analytic nonadiabatic derivative couplings, a feature that is currently under development. However, due to slight spin contamination, nonadiabatic couplings between states of different multiplicities may be nonzero, and fully avoiding such unphysical couplings relies on appropriate spin-adaptation within the noncollinear spin-flip TDDFT framework. Moreover, one notable advantage of the noncollinear spin-flip TDDFT approach is that it offers a consistent framework for treating both derivative couplings and spin-orbit coupling effects simultaneously, which is particularly beneficial for systems where intersystem crossing and internal conversion are energetically competitivemarian2012spin.
@@
    Description of single-bond dissociation. Nonadiabatic dynamics simulations may involve significant changes in bond lengths and angles, and even chemical reactions, in which case strong multiconfigurational character may arise. Noncollinear spin-flip TDDFT is particularly well-suited to describe such situations, as it allows the simultaneous treatment of several important configurations, including $|{\mathrm{CSS}}\rangle$, $|{\mathrm{OSS}}\rangle$, $|{\mathrm{DES}}\rangle$. In Appendix C, noncollinear spin-flip TDDFT is utilized to investigate single-bond dissociation and is compared with broken-symmetry unrestricted Kohn–Sham (UKS) DFT.
@@
    Computational efficiency. Computational cost is always a key consideration for practical dynamics simulations. As discussed in Ref. wang2025zero, the computational cost of noncollinear spin-flip TDDFT and its gradients is comparable to that of DFT and spin-conserving TDDFT, significantly lower than that of traditional multireference wavefunction methods. In this work, the computational timings of noncollinear spin-flip TDDFT is further compared with collinear spin-flip TDDFT, with the results shown in Table S5 in the Supporting Information. The noncollinear version is indeed more computationally demanding, but the difference diminishes as the system size increases.
5 Conclusions and Outlook
Noncollinear spin-flip TDDFT, like conventional spin-conserving TDDFT, has a rigorous theoretical foundation. With multicollinear functionals, it can correctly capture the topological features near potential energy surface crossings and provide reasonable spin multiplicities without extra approximations. Compared to traditional DFT/TDDFT methods, its advantage lies in successfully handling $\mathrm{S}_{0}/\mathrm{S}_{1}$ conical intersections. Compared to collinear spin-flip TDDFT, it has three main advantages: First, it significantly reduces spin contamination near conical intersections. Second, it can accurately describe singlet-triplet crossings without the self-splitting issue. Third, it does not impose specific requirements on the choice of functional, free of artificial potential energy surface crossings caused by degenerate orbitals even for pure functionals.
While the method encounters certain limitations, such as slight spin contamination in low-lying excited states and inadequate spin description for some high-lying states, its low computational cost still makes it a promising electronic structure tool for nonadiabatic dynamics simulations, especially for larger systems.
Appendix
Appendix A Basis Set and Functional Dependence of MECI Geometries
To assess the dependence of MECI geometries on the basis set, we optimize the MECI of ethylene using noncollinear spin-flip TDDFT with BHHLYP functional and a series of basis setsditchfield1971selfconsistent, hehre1972self, hariharan1973influence, clark1983efficient, dunning1989gaussian, kendall1992electron. The results are summarized in Table 4. Using the largest basis set, aug-cc-pVQZ, as the reference, RMSDs of the optimized geometries remain very small–around or below $0.01\text{\,}\mathrm{\SIUnitSymbolAngstrom}$–indicating that the choice of basis set has a negligible effect on the computed MECI geometries.
Table 4: RMSD of MECI geometries of MECI geometries of ethylene obtained by noncollinear spin-flip TDDFT (BHHLYP) using different basis sets relative to those computed using the aug-$c$-pVQZ basis set.
Basis Set Number of Basis Functions RMSD ($\mathrm{\SIUnitSymbolAngstrom}$)
6-31G* 36 0.014
6-31+G** 56 0.008
$c$-pVDZ 48 0.011
$c$-pVTZ 116 0.003
$c$-pVQZ 230 0.002
aug-$c$-pVDZ 82 0.011
aug-$c$-pVTZ 184 0.000
aug-$c$-pVQZ 344
The effect of the exchange-correlation functional is also examined by optimizing the MECIs of ethylene, PSB3, and styrene using noncollinear spin-flip TDDFT with the 6-31+G** basis set and different functionalsdirac1930note, vosko1980accurate, becke1988densityfunctional, lee1988development, becke1993densityfunctional, stephens1994initio, perdew1996generalized, adamo1999reliable, yanai2004new, tao2003climbing, zhao2008m06. The resulting geometries are compared with high-level MRCISD datanikiforov2014assessment and the corresponding RMSDs are summarized in Table 5. In contrast to the basis-set dependence, the choice of functional shows a more pronounced influence on the MECI geometries. Hybrid functionals generally perform better than pure functionals in reproducing the MECI geometries, with BHHLYP giving the best results among the tested functionals for the present systems. On the other hand, pure functionals tend to yield smaller spin contamination.
Table 5: RMSD (in $\mathrm{\SIUnitSymbolAngstrom}$) of MECI geometries obtained by noncollinear spin-flip TDDFT with various functionals relative to MRCISD resultsnikiforov2014assessment and $\langle{\hat{S}^{2}}\rangle$ values of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ states at the MECI geometries.
Functionals Ethylene PSB3 Styrene
RMSD $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{0}}$ $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{1}}$ RMSD $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{0}}$ $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{1}}$ RMSD $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{0}}$ $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{1}}$
SVWN 0.215 0.008 0.014 0.490 0.009 0.013 0.145 0.011 0.015
BLYP 0.201 0.009 0.026 0.375 0.035 0.024 0.126 0.020 0.027
B3LYP 0.103 0.012 0.050 0.248 0.027 0.050 0.104 0.043 0.046
BHHLYP 0.018 0.023 0.092 0.035 0.071 0.107 0.031 0.113 0.117
PBE 0.188 0.011 0.029 0.335 0.033 0.028 0.116 0.022 0.029
PBE0 0.028 0.017 0.060 0.169 0.037 0.068 0.099 0.057 0.061
CAM-B3LYP 0.051 0.014 0.063 0.032 0.071 0.072 0.101 0.068 0.076
TPSS 0.106 0.014 0.043 0.365 0.064 0.041 0.107 0.034 0.040
M06-2X 0.118 0.016 0.105 0.057 0.100 0.054 0.113 0.037 0.108
Appendix B Calculations of Excited States Dominated by Doubly Excited Configuration